40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4885     Entry time: Sun Jun 26 16:02:12 2011
Author: kiwamu 
Type: Update 
Category: IOO 
Subject: Friday MC activity 

[Rana / Kiwamu]

 Last Friday we did several things for MC :

   - aligned the incident beam to MC

   - increased the locking gain by 6 dB and modified the auto-locker script accordingly

   - improved the alignment of the beam on the MC_REFLPD photo diode

 


(Motivation)

 In the beginning of the work, we wanted to know what RF frequency components are prominent in the reflection from MC.

Since the WFS circuits are capable for two RF notches, we wanted to determine which frequencies are appropriate for those notches.

So for the purpose we tried searching for unwanted RF components in the reflection.

However during the work, we found several things that needed to be fixed, so we spent most of the time for improving the MC locking.

 

(Some notes)

 - Alignments of the incident beam

At the beginning, the reflection from MC was about 2.2 in C1:IOO-REFLDC and the lock of MC had been frequently unlocked.

This situation of high reflection seemed to be related to a work done by Suresh (#4880).

Rana went to the PSL table and tweaked two input steering mirrors in the zig-zag path, and finally the reflection went down to ~ 0.8 in C1:IOO-REFLDC.

This work made the lock more robust.

 

 - Change of the locking gain

 After the alignment of the incident beam, we started looking at the time series of the MC_REFLPD signal with an oscilloscope as a start point.

What we found was a significant amount of 30 kHz components. This 30 kHz oscillation was thought be a loop oscillation, and indeed it was so.

We increased the loop gain by 6 dB and then the 30 kHz components disappeared successfully.

So the nominal locking gain of MC is now 11 dB in C1:IOO-MC_REFL_GAIN. The auto locker script was also modified accordingly.

 

- RF components in the MCREFL signal

After those improvements mentioned above, we started looking at the spectrum of the MCREFL PD using the spectrum analyzer HP8590.

The 29.5 MHz component was the biggest components in the spectrum. Ideally this 29.5 MHz signal should be zero when MC is locked.

One possible reason for this big 29.5 MHz signal was because the lock point was off from the resonant point.

We tweaked the offset in the MC lock path using a digital offset, C1:IOO-MC-REFL_OFFSET.

We found an offset point where the 29.5MHz signal went to the minimum, but didn't go to zero.

 

(works to be done)

So it needs some more works to investigate the cause of nonzero 29.5 MHz signal as well as investigation of what RF components should be notched out.

A good start point would be writing a GPIB interface script such that we can get the spectra from HP8590 without any pains.

ELOG V3.1.3-